### v1

# Generalized mKdV Equation and Genus Two Jacobi Type Hyperelliptic Differential Equation

Generalized mKdV Equation and Genus Two Jacobi Type Hyperelliptic Differential Equation
Masahito Hayashi
Osaka Institute of Technology, Osaka 535-8585, Japan
Kazuyasu Shigemoto
Tezukayama University, Nara 631-8501, Japan
Takuya Tsukioka
Bukkyo University, Kyoto 603-8301, Japan
masahito.hayashi@oit.ac.jpshigemot@tezukayama-u.ac.jptsukioka@bukkyo-u.ac.jp
We generalized the mKdV equation in order that the static equations include $\rm sn$} differential equation. As a result, a good correspondence was obtained between the KdV equation and the mKdV equation. For general genus two hyperelliptic curves, we obtained differential equations for Weierstrass type and Jacobi type hyperelliptic functions. Considering the special case of $\lambda_{6}=0,\lambda_{0}=0$, Weierstrass type and Jacobi type hyperelliptic functions are different solutions to the same hyperelliptic differential equations. Then these solutions are connected by the special $\rm Sp(4,{\bf R})$} Lie group transformation.
Keywords: generalized mKdV equation, genus two, Jacobi type, hyperelliptic differential equations
1 Introduction
The reason why soliton system can be solved exactly despite being nonlinear is that there is the Lie group structure. It is well known that the KdV equation has the ${\rm Sp(2,{\bf R})/Z_{2}}\cong{\rm SO(2,1)}$ Lie group structure [1, 2]. We have demonstrated that $\wp$ function, which is the special static solution of the KdV equation, has the ${\rm Sp(2,{\bf R})/Z_{2}}\cong{\rm SO(2,1)}$ Lie group structure [3], and the genus two hyperelliptic $\wp$ functions have the ${\rm Sp(4,{\bf R})/Z_{2}}\cong{\rm SO(3,2)}$ Lie group structure [3, 4, 5]. Not in the general Riemann surfaces but in the hyperelliptic curves, we expect that multi-periodic functions, i.e., hyperelliptic functions have the Lie group structure.
In this paper, we first generalized mKdV equation in order that the static differential equation includes $\rm sn$} differential equation. This gives a good correspondence between the KdV equation and the mKdV equation. For general genus two hyperelliptic curves, we obtain the conventional differential equations for Weierstrass type and Jacobi type hyperelliptic functions. If we consider the special case of $\lambda_{6}=0,\lambda_{0}=0$, differential equations for Weierstrass type and Jacobi type hyperelliptic functions are different solutions of the same hyperelliptic differential equations. We confirmed that two hyperelliptic functions are linked by a special half-period $\rm Sp(4,{\bf R})$} Lie group transformation.
2 Generalized mKdV equation
2.1 Integrability condition for the generalized mKdV equation
The KdV equation is expressed as follows
$u_{t}(x,t)+u_{xxx}(x,t)-6u(x,t)u_{x}(x,t)=0,$ (2.1)
and mKdV equation is expressed as follows
$v_{t}(x,t)+v_{xxx}(x,t)-6v^{2}(x,t)v_{x}(x,t)=0,$ (2.2)
where we denote $\dfrac{\partial u(x,t)}{\partial t}=u_{t}(x,t)$, $\dfrac{\partial u(x,t)}{\partial x}=u_{x}(x,t)$, $\dfrac{\partial^{3}u(x,t)}{\partial x^{3}}=u_{xxx}(x,t)\rule$[-9.95845pt]{0.0pt}{25.60747pt} etc. Eq.(2.2) is generalized in order that the static equation includes differential equation for the $\rm sn$} function as follows
$v_{t}(x,t)+v_{xxx}(x,t)-6v^{2}(x,t)v_{x}(x,t)+av_{x}(x,t)=0.$ (2.3)
We call Eq.(2.3) as the generalized mKdV equation. We demonstrate that this generalized mKdV equation is an integrable system. To achieve this, we adopt the following AKNS formalism [7, 6]
$\frac{\partial}{\partial x}\left(\begin{array}[]{c}\psi_{1}\\ \psi_{2}\end{array}\right)=\left(\begin{array}[]{cc}\eta&v(x,t)\\ -v(x,t)&-\eta\end{array}\right)\left(\begin{array}[]{c}\psi_{1}\\ \psi_{2}\end{array}\right)=L\left(\begin{array}[]{c}\psi_{1}\\ \psi_{2}\end{array}\right),$ (2.12)
$\frac{\partial}{\partial t}\left(\begin{array}[]{c}\psi_{1}\\ \psi_{2}\end{array}\right)=\left(\begin{array}[]{cc}A&B\\ C&-A\end{array}\right)\left(\begin{array}[]{c}\psi_{1}\\ \psi_{2}\end{array}\right)=M\left(\begin{array}[]{c}\psi_{1}\\ \psi_{2}\end{array}\right),$ (2.21)
$A=2\eta v^{2}(x,t)-4\eta^{3}b,$ (2.22)
$B=-v_{xx}(x,t)-2\eta v_{x}(x,t)+2v^{3}(x,t)-4\eta^{2}bv(x,t),$ (2.23)
$C=v_{xx}(x,t)+6\eta v_{x}(x,t)-2v^{3}(x,t)+4\eta^{2}bv(x,t).$ (2.24)
$\eta$ is the spectral parameter and $b$ is a constant parameter, which is determined later. The matrix $M$ in Eq.(2.21) has $\rm GL(2,{\bf R})$} structure. The integrability condition of Eqs.(2.1) and (2.2) is given by
$\left[\frac{\partial}{\partial x}-L,\frac{\partial}{\partial t}-M\right]=\frac{\partial L}{\partial t}-\frac{\partial M}{\partial x}+\left[L,M\right]=\left(\begin{array}[]{cc}0&D\\ -D&0\end{array}\right)=0,$ (2.27)
$D=v_{t}(x,t)+v_{xxx}(x,t)-6v^{2}(x,t)v_{x}(x,t)+4\eta^{2}(b-1)v_{x}(x,t),$ (2.28)
and $b$ is set to be $a=4\eta^{2}(b-1)$. Then the above integrability condition $D=0$ gives the generalized mKdV equation.
2.2 Various transformations between the static KdV and the static generalized mKdV equations
For simplicity, we will denote $v(x,0)$ as $v(x)$. This type of abbreviation is used in other cases as well.
2.2.1 Miura transformation
Using the static Miura transformation $u(x)=v^{2}(x)+v_{x}(x)-a/6$, we obtain
$u_{xxx}(x)-6u(x)u_{x}(x)=\left(\frac{d}{dx}+2v(x)\right)\left(v_{xxx}(x)-6v^{2}(x)v_{x}(x)+av_{x}(x)\right).$ (2.29)
Therefore, if
$v_{xxx}(x)-6v^{2}(x)v_{x}(x)+av_{x}(x)=0,$ (2.30)
we obtain the static KdV equation
$u_{xxx}(x)-6u(x)u_{x}(x)=0.$ (2.31)
Of course, the time dependent KdV and generalized mKdV equations are linked by the time dependent Miura transformation $u(x,t)=v^{2}(x,t)+v_{x}(x,t)-a/6$ [8].
2.2.2 Square transformation
Using the square transformation $u(x)=2v^{2}(x)-2a/3$, we obtain
$u_{xxx}(x)-6u(x)u_{x}(x)=\left(4v(x)\frac{d}{dx}+12v_{x}(x)\right)\left(v_{xx}(x)-2v^{3}(x)+av(x)\right).$ (2.32)
Therefore, if
$v_{xx}(x)-2v^{3}(x)+av(x)=0,$ (2.33)
we obtain the static KdV equation
$u_{xxx}(x)-6u(x)u_{x}(x)=0.$ (2.34)
Note that
$\frac{d}{dx}\left(v_{xx}(x)-2v^{3}(x)+av(x)\right)=v_{xxx}(x)-6v^{2}(x)v_{x}(x)+av_{x}(x).$ (2.35)
Eq.(2.33) means this $v(x)$ satisfies the static generalized mKdV equation.
2.2.3 Inverse square transformation
Using the inverse square transformation $u(x)=1/v^{2}(x)-2a/3$, we obtain
$u_{xxx}$ $(x)-6u(x)u_{x}(x)$
$=$ $\left(-\frac{1}{v^{3}(x)v_{x}(x)}\frac{d^{2}}{dx^{2}}+\left(\frac{v_{xx}(x)}{v^{3}(x){v_{x}^{2}(x)}}+\frac{9}{v^{4}(x)}\right)\frac{d}{dx}-\frac{24v_{x}(x)}{v^{5}(x)}\right)$
$\times\left(v_{x}^{2}(x)-v^{4}(x)+av^{2}(x)-\frac{1}{2}\right).$ (2.36)
Therefore, if
$v_{x}^{2}(x)-v^{4}(x)+av^{2}(x)-\frac{1}{2}=0,$ (2.37)
we obtain the static KdV equation
$u_{xxx}(x)-6u(x)u_{x}(x)=0.$
Note that
$\frac{d}{dx}\left(v_{x}^{2}(x)-v^{4}(x)+av^{2}(x)-\frac{1}{2}\right)=2v_{x}(x)\left(v_{xx}(x)-2v^{3}(x)+av(x)\right).$ (2.38)
In this case, $v(x)$ not only satisfies the static generalized mKdV equation but also becomes the Jacobi $\rm sn$} function. This transformation is well-known that links the Weierstrass $\wp$ function and the Jacobi $\rm sn$} function in the form
$\wp\left(\frac{u}{\sqrt{e_{1}-e_{3}}}\right)=e_{3}+\frac{(e_{1}-e_{3})}{{\rm sn}^{2}(u)}.$
2.2.4 Inverse power transformation
Using the inverse power transformation u(x)=1/v(x)a/6u(x)=1/v(x)-a/6, we obtain
$u_{xxx}$ $(x)-6u(x)u_{x}(x)$
$=$ $\left(-\frac{1}{2v^{2}(x)v_{x}(x)}\frac{d^{2}}{dx^{2}}+\left(\frac{v_{xx}(x)}{2v^{2}(x){v_{x}(x)}^{2}}+\frac{3}{v^{3}(x)}\right)\frac{d}{dx}-\frac{6v_{x}(x)}{v^{4}(x)}\right)$
$\times\left(v_{x}^{2}(x)-v^{4}(x)+av^{2}(x)-2v(x)\right).$ (2.39)
Therefore, if
$v_{x}^{2}(x)-v^{4}(x)+av^{2}(x)-2v(x)=0,$ (2.40)
we obtain the static KdV equation
$u_{xxx}(x)-6u(x)u_{x}(x)=0.$
Note that
$\frac{d}{dx}\left(v_{x}^{2}(x)-v^{4}(x)+av^{2}(x)-2v(x)\right)=2v_{x}(x)\left(v_{xx}(x)-2v^{3}(x)+av(x)-1\right),$
and
$\frac{d}{dx}\left(v_{xx}(x)-2v^{3}(x)+av(x)-1\right)=v_{xxx}(x)-6v^{2}(x)v_{x}(x)+av_{x}(x).$
In this case, $v(x)$ is a solution of the static generalized mKdV equation, but it is not that of the Jacobi $\rm sn$} function.
2.2.5 A relationship between transformations
Comparing the square transformation and the inverse square transformation, we obtain
$u(x)=2{v_{1}^{2}(x)}-\frac{2a}{3}=\frac{1}{{v_{2}^{2}(x)}}-\frac{2a}{3},$ (2.41)
where $v_{1}(x)$ and $v_{2}(x)$ satisfies the same $\rm sn$} type differential equation. In order that Eq.(2.40) becomes the standard $\rm sn$} differential equation, we put $k=\sqrt{2}$, $a=(1+k^{2})/2=3/2$, and we obtain [3]
$v_{1}(x)={\rm sn}(x/\sqrt{2})~~\mathrm{and}~~v_{2}(x)=\widehat{{\rm sn}}(x/\sqrt{2})=\frac{1}{\sqrt{2}\ {\rm sn}(x/\sqrt{2})}=\frac{1}{\sqrt{2}\ v_{1}(x)}.$ (2.42)
More generally, using ${\rm sn}_{z}^{2}(z)=\left(1-{\rm sn}^{2}(z)\right)\left(1-k^{2}{\rm sn}^{2}(z)\right)$, the differential equation for ${\rm sn}(z)$ becomes
${\rm sn}_{zz}(z)=-\left(1+k^{2}\right){\rm sn}(z)+2k^{2}{\rm sn}^{3}(z).$ (2.43)
On the other hand, ${\widehat{{\rm sn}}(z)=\frac{1}{k\,{\rm sn}(z)}}$ satisfies the same differential equation
$\widehat{{\rm sn}}_{zz}(z)=-\left(1+k^{2}\right)\widehat{{\rm sn}}(z)+2k^{2}\widehat{{\rm sn}}^{3}(z).$ (2.44)
The reason for this relation comes from the half-period relation ${{\rm sn}(z+3iK^{\prime})=\frac{1}{k\ {\rm sn}(z)}}$.
3 Differential equations for genus two Jacobi type hyperelliptic functions
We start from the genus two hyperelliptic curves of the form
${y_{i}}^{2}=\lambda_{6}{x_{i}}^{6}+\lambda_{5}{x_{i}}^{5}+\lambda_{4}{x_{i}}^{4}+\lambda_{3}{x_{i}}^{3}+\lambda_{2}{x_{i}}^{2}+\lambda_{1}{x_{i}}+\lambda_{0},\quad(i=1,2).$ (3.1)
The Jacobi inversion relation is given by
$du_{1}=\frac{dx_{1}}{y_{1}}+\frac{dx_{2}}{y_{2}},\quad du_{2}=\frac{x_{1}dx_{1}}{y_{1}}+\frac{x_{2}dx_{2}}{y_{2}},$ (3.2)
which gives
$\frac{\partial x_{1}}{\partial u_{2}}=\frac{y_{1}}{x_{1}-x_{2}},\quad\frac{\partial x_{2}}{\partial u_{2}}=-\frac{y_{2}}{x_{1}-x_{2}},\quad\frac{\partial x_{1}}{\partial u_{1}}=-\frac{x_{2}y_{1}}{x_{1}-x_{2}},\quad\frac{\partial x_{2}}{\partial u_{1}}=\frac{x_{1}y_{2}}{x_{1}-x_{2}}.$ (3.3)
We define
$F(x_{1},x_{2};\left\{\lambda_{j}\right\})=$ $2\lambda_{6}x_{1}^{3}x_{2}^{3}+\lambda_{5}x_{1}^{2}x_{2}^{2}(x_{1}+x_{2})+2\lambda_{4}x_{1}^{2}x_{2}^{2}+\lambda_{3}x_{1}x_{2}(x_{1}+x_{2})$
$+2\lambda_{2}x_{1}x_{2}+\lambda_{1}(x_{1}+x_{2})+2\lambda_{0}.$ (3.4)
Then the genus two Weierstrass type hyperelliptic functions are defined as follows [9]
$\Re_{22}(u_{1},u_{2})=\frac{\lambda_{5}}{4}\left(x_{1}+x_{2}+\frac{2\lambda_{6}}{\lambda_{5}}(x_{1}^{2}+x_{1}x_{2}+x_{2}^{2})\right),$ (3.5)
$\Re_{21}(u_{1},u_{2})=\frac{\lambda_{5}}{4}\left(-x_{1}x_{2}-\frac{2\lambda_{6}}{\lambda_{5}}x_{1}x_{2}(x_{1}+x_{2})\right),$ (3.6)
$\Re_{11}(u_{1},u_{2})=\frac{F(x_{1},x_{2};\left\{\lambda_{j}\right\})-2y_{1}y_{2}}{4(x_{1}-x_{2})^{2}}+\frac{\lambda_{6}}{2}x_{1}^{2}x_{2}^{2}.$ (3.7)
[…]
$\frac{\partial\Re_{22}}{\partial u_{1}}=\frac{\partial\Re_{21}}{\partial u_{2}},\quad\frac{\partial\Re_{21}}{\partial u_{1}}=\frac{\partial\Re_{11}}{\partial u_{2}}.$ (3.8)
Though $\Re_{22}$ and $\Re_{21}$ are symmetric functions, but they are not the fundamental symmetric functions. Then any symmetric functions cannot be expressed in the closed form by $\Re_{22}$ and $\Re_{21}$. So far, closed differential equations for the genus two hyperelliptic functions are known only for the $\lambda_{6}=0$ case [10].
3.1 Genus two Weierstrass type hyperelliptic differential equations
Even for the general genus two hyperelliptic curves with $\lambda_{6}\neq 0$, we consider the genus two hyperelliptic functions $\wp_{22}$, $\wp_{21}$ and $Q$ of the form
$\wp_{22}(u_{1},u_{2})=\frac{\lambda_{5}}{4}(x_{1}+x_{2}),\ \wp_{21}(u_{1},u_{2})=-\frac{\lambda_{5}}{4}x_{1}x_{2},\ Q(u_{1},u_{2})=\frac{F(x_{1},x_{2};\left\{\lambda_{j}\right\})-2y_{1}y_{2}}{4(x_{1}-x_{2})^{2}}.$ (3.9)
We call the above as the Weierstrass type hyperelliptic functions. Here $\wp_{22}$ and $\wp_{21}$ are fundamental symmetric function. Then any symmetric quantities can be expressed by $\wp_{22}$ and $\wp_{21}$. Thus we can obtain the hyperelliptic differential equations in the closed forms. Though we obtain the relation
$\frac{\partial\wp_{22}}{\partial u_{1}}=\frac{\partial\wp_{21}}{\partial u_{2}},\quad\frac{\partial\wp_{21}}{\partial u_{1}}\neq\frac{\partial Q}{\partial u_{2}},$ (3.10)
the condition ${\frac{\partial\wp_{22}}{\partial u_{1}}=\frac{\partial\wp_{21}}{\partial u_{2}}}$ alone is sufficient to guarantee the integrability condition, i.e. the existence of $\sigma(u_{1},u_{2})$ function as a potential. If $\wp_{22}$ and $\wp_{21}$ are given as functions of $\sigma$ function, $Q$ can in principle be given as a function of $\sigma$ function via the generalized Kummer surface relation. We will give the generalized Kummer surface relation later.
In order to obtain the differential equations for genus two Weierstrass type hyperelliptic functions, we first calculate $\wp_{2222}$, $\wp_{2221}$, $\wp_{2211}$, $\wp_{2111}$, $Q_{11}$, $Q_{21}$ and $Q_{22}$. Because these quantities are symmetric functions of $x_{1}$ and $x_{2}$, these are expressed by $\wp_{22}$ and $\wp_{21}$. If these quantities contain symmetric function of the type $y_{1}y_{2}$, the quantity $y_{1}y_{2}$ is expressed by $Q$ and/or $Q^{2}$ by using ${y_{1}y_{2}=-2(x_{1}-x_{2})^{2}Q+\frac{1}{2}F(x_{1},x_{2})}$,
Then we obtain the differential equations of the genus two Weierstrass type hyperelliptic functions of the form
$1)\ \wp_{2222}-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{3}-\frac{12\lambda_{6}}{\lambda_{5}}\wp_{22}\wp_{21}-6\wp_{22}^{2}-\lambda_{4}\wp_{22}-\lambda_{5}\wp_{21}-\frac{\lambda_{3}\lambda_{5}}{8}=0,$ (3.11)
$2)\ \wp_{2221}-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{2}\wp_{21}-\frac{4\lambda_{6}}{\lambda_{5}}\wp_{21}^{2}-6\wp_{22}\wp_{21}-\lambda_{4}\wp_{21}+\frac{\lambda_{5}}{2}Q=0,$ (3.12)
$3)\ \wp_{2211}-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}\wp_{21}^{2}-2\wp_{22}Q-4\wp_{21}^{2}-\frac{\lambda_{3}}{2}\wp_{21}=0,$ (3.13)
$4)\ \wp_{2111}-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{21}^{3}-6\wp_{21}Q+\frac{\lambda_{1}}{2}\wp_{22}-\lambda_{2}\wp_{21}+\frac{\lambda_{0}\lambda_{5}}{4}=0,$ (3.14)
$5)\ Q_{11}-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{21}^{2}Q+\frac{16\lambda_{0}\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{2}-\frac{8\lambda_{1}\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}\wp_{21}-6Q^{2}+3\lambda_{0}\wp_{22}$
$-\left(\lambda_{1}-\frac{2\lambda_{0}\lambda_{6}}{\lambda_{5}}\right)\wp_{21}-\lambda_{2}Q+\frac{\lambda_{0}\lambda_{4}}{2}-\frac{\lambda_{1}\lambda_{3}}{8}=0,$ (3.15)
$6)\ Q_{21}-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}\wp_{21}Q+\frac{4\lambda_{1}\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{2}-\frac{8\lambda_{2}\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}\wp_{21}+\frac{4\lambda_{3}\lambda_{6}}{\lambda_{5}^{2}}\wp_{21}^{2}$
$-6\wp_{21}Q+\left(\frac{2\lambda_{0}\lambda_{6}}{\lambda_{5}}+\frac{\lambda_{1}}{2}\right)\wp_{22}-\lambda_{2}\wp_{21}+\frac{\lambda_{0}\lambda_{5}}{4}=0,$ (3.16)
$7)\ Q_{22}-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{2}Q+\left(\frac{16\lambda_{4}\lambda_{6}}{\lambda_{5}^{2}}-4\right)\wp_{21}^{2}-\frac{8\lambda_{3}\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}\wp_{21}-2\wp_{22}Q$
$-\frac{20\lambda_{6}}{\lambda_{5}}\wp_{21}Q+\frac{\lambda_{1}\lambda_{6}}{\lambda_{5}}\wp_{22}+\left(-\frac{2\lambda_{2}\lambda_{6}}{\lambda_{5}}-\frac{\lambda_{3}}{2}\right)\wp_{21}+\frac{\lambda_{0}\lambda_{6}}{2}=0.$ (3.17)
From Eq.(3.12), we naturally obtain $Q$ in the combination of Eq.(3.9).
The characteristic feature of $\lambda_{6}\neq 0$ hyperelliptic differential equations is the existence of the cubic term, that is, terms ${-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{3},-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{2}\wp_{21},\cdots,-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{2}Q}$, in Eq.(3.11) $\sim$ Eq.(3.17).
Here we give the generalized Kummer surface relation. For $\wp_{22}$, $\wp_{21}$ and $Q$, only two are independent. Then there exist one relation, which is called the generalized Kummer surface relation. By using the relation ${y_{1}y_{2}=-2(x_{1}-x_{2})^{2}Q+\frac{1}{2}F(x_{1},x_{2})}$, we obtain
${y_{1}^{2}y_{2}^{2}=4(x_{1}-x_{2})^{4}Q^{2}-2(x_{1}-x_{2})^{2}QF(x_{1},x_{2})+\frac{1}{4}F(x_{1},x_{2})^{2}}.$
Using Eq.(3.1), this constraint is expressed by $\wp_{22}$, $\wp_{21}$ and $Q$. The explicit expression is given as follows. We denote $X=\wp_{22}$, $Y=\wp_{21}$, $Z=Q$, and we obtain the following generalized Kummer surface relation of the form
$K_{2}=K_{1}+\frac{4\lambda_{6}}{\lambda_{5}^{2}}\Big(-\lambda_{5}^{2}Y^{2}-8\lambda_{0}\lambda_{5}X^{2}Y+4\lambda_{1}\lambda_{5}XY^{2}$
$+16(-\lambda_{0}X^{4}+\lambda_{1}X^{3}Y-\lambda_{2}X^{2}Y^{2}+\lambda_{3}XY^{3}-\lambda_{4}Y^{4}+\lambda_{5}Y^{3}Z)\Big)=0,$ (3.18)
$K_{1}=\left|\begin{array}[]{cccc}-\lambda_{0}&\lambda_{1}/2&2Z&-2Y\\ \lambda_{1}/2&-\lambda_{2}-4Z&\lambda_{3}/2+2Y&2X\\ 2Z&\lambda_{3}/2+2Y&-\lambda_{4}-4X&\lambda_{5}/2\\ -2Y&2X&\lambda_{5}/2&0\\ \end{array}\right|.$ (3.23)
In the $\lambda_{6}=0$ case, the generalized Kummer surface relation $K_{2}=0$ reduce to the standard Kummer surface relation $K_{1}=0$ [11].
If we consider the special case of $\lambda_{6}=0$, we obtain ${\frac{\partial\wp_{21}}{\partial u_{1}}=\frac{\partial Q}{\partial u_{2}}}$, then we can put $Q=\wp_{11}$. Thus we obtain the conventional differential equations for Weierstrass type hyperelliptic functions [11]. If we consider the more special case of $\lambda_{6}=0,\lambda_{0}=0$, differential equations for Weierstrass type hyperelliptic differential functions become
$1)\ \wp_{2222}-6\wp_{22}^{2}-\lambda_{4}\wp_{22}-\lambda_{5}\wp_{21}-\frac{\lambda_{3}\lambda_{5}}{8}=0,$ (3.24)
$2)\ \wp_{2221}-6\wp_{22}\wp_{21}-\lambda_{4}\wp_{21}+\frac{\lambda_{5}}{2}\wp_{11}=0,$ (3.25)
$3)\ \wp_{2211}-2\wp_{22}\wp_{11}-4\wp_{21}^{2}-\frac{\lambda_{3}}{2}\wp_{21}=0,$ (3.26)
$4)\ \wp_{2111}-6\wp_{21}\wp_{11}+\frac{\lambda_{1}}{2}\wp_{22}-\lambda_{2}\wp_{21}=0,$ (3.27)
$5)\ \wp_{1111}-6\wp_{11}^{2}-\lambda_{1}\wp_{21}-\lambda_{2}\wp_{11}-\frac{\lambda_{1}\lambda_{3}}{8}=0.$ (3.28)
The condition $\lambda_{0}=0$ is not the special restriction, because we can always put $\lambda_{0}=0$ by the constant shift of $x_{i}$ in the hyperelliptic curve.
3.2 Genus two Jacobi type hyperelliptic differential equations
Next, we try to obtain differential equations for the genus two Jacobi type hyperelliptic functions by the following dual transformation. Under the dual transformation
$x_{i}\rightarrow\frac{1}{x_{i}},\quad y_{i}\rightarrow\frac{y_{i}}{x_{i}^{3}},\quad\lambda_{j}\leftrightarrow\lambda_{6-j},\ (j=0,1,\cdots,6),$
hyperelliptic curves of Eq.(3.1) are invariant. Under this transformation, we obtain
$du_{1}=\frac{dx_{1}}{y_{1}}+\frac{dx_{2}}{y_{2}}\rightarrow-\left(\frac{x_{1}dx_{1}}{y_{1}}+\frac{x_{2}dx_{2}}{y_{2}}\right)=-du_{2},$ (3.29)
$du_{2}=\frac{x_{1}dx_{1}}{y_{1}}+\frac{x_{2}dx_{2}}{y_{2}}\rightarrow-\left(\frac{dx_{1}}{y_{1}}+\frac{dx_{2}}{y_{2}}\right)=-du_{1}.$ (3.30)
Then the corresponding hyperelliptic functions become of the form
$\wp_{22}(u_{1},u_{2})=\frac{\lambda_{5}}{4}(x_{1}+x_{2})\rightarrow\widehat{\wp}_{11}(u_{1},u_{2})=\frac{\lambda_{1}}{4}\left(\frac{1}{x_{1}}+\frac{1}{x_{2}}\right),$ (3.31)
$\wp_{21}(u_{1},u_{2})=\frac{\lambda_{5}}{4}x_{1}x_{2}\rightarrow\widehat{\wp}_{21}(u_{1},u_{2})=-\frac{\lambda_{1}}{4}\frac{1}{x_{1}x_{2}},$ (3.32)
$Q(u_{1},u_{2})\rightarrow\ \widehat{Q}(u_{1},u_{2})=\frac{F(x_{1},x_{2};\left\{\lambda_{j}\right\})-2y_{1}y_{2}}{4x_{1}x_{2}(x_{1}-x_{2})^{2}}.$ (3.33)
We call these hyperelliptic functions $\widehat{\wp}_{11}$, $\widehat{\wp}_{21}$ and $\widehat{Q}$ as the Jacobi type. These satisfy relations
$\frac{\partial\widehat{Q}}{\partial u_{1}}\neq\frac{\partial\widehat{\wp}_{21}}{\partial u_{2}},\quad\frac{\partial\widehat{\wp}_{21}}{\partial u_{1}}=\frac{\partial\widehat{\wp}_{11}}{\partial u_{2}}.$ (3.34)
Differential equations for the genus two Jacobi type hyperelliptic functions are given by
$1)\ \widehat{\wp}_{2221}-6\widehat{\wp}_{21}\widehat{Q}-\lambda_{4}\widehat{\wp}_{21}+\frac{\lambda_{5}}{2}\widehat{\wp}_{11}-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{3}+\frac{\lambda_{1}\lambda_{6}}{4}=0,$ (3.35)
$2)\ \widehat{\wp}_{2211}-2\widehat{\wp}_{11}\widehat{Q}-4\widehat{\wp}_{21}^{2}-\frac{\lambda_{3}}{2}\widehat{\wp}_{21}-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{2}\widehat{\wp}_{11}=0,$ (3.36)
$3)\ \widehat{\wp}_{2111}-6\widehat{\wp}_{21}\widehat{\wp}_{11}+\frac{\lambda_{1}}{2}\widehat{Q}-\lambda_{2}\widehat{\wp}_{21}-\frac{4\lambda_{0}}{\lambda_{1}}\widehat{\wp}_{21}^{2}-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}\widehat{\wp}_{11}^{2}=0,$ (3.37)
$4)\ \widehat{\wp}_{1111}-6\widehat{\wp}_{11}^{2}-\lambda_{1}\widehat{\wp}_{21}-\lambda_{2}\widehat{\wp}_{11}-\frac{12\lambda_{0}}{\lambda_{1}}\widehat{\wp}_{21}\widehat{\wp}_{11}-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{11}^{3}-\frac{\lambda_{1}\lambda_{3}}{8}=0,$ (3.38)
$5)\ \widehat{Q}_{11}-2\widehat{\wp}_{11}\widehat{Q}-\left(4-\frac{16\lambda_{0}\lambda_{2}}{\lambda_{1}^{2}}\right)\widehat{\wp}_{21}^{2}+\frac{\lambda_{0}\lambda_{5}}{\lambda_{1}}\widehat{\wp}_{11}-\left(\frac{\lambda_{3}}{2}+\frac{2\lambda_{0}\lambda_{4}}{\lambda_{1}}\right)\widehat{\wp}_{21}$
$-\frac{20\lambda_{0}}{\lambda_{1}}\widehat{\wp}_{21}\widehat{Q}-\frac{8\lambda_{0}\lambda_{3}}{\lambda_{1}^{2}}\widehat{\wp}_{21}\widehat{\wp}_{11}-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{11}^{2}\widehat{Q}+\frac{\lambda_{0}\lambda_{6}}{2}=0,$ (3.39)
$6)\ \widehat{Q}_{21}-6\widehat{\wp}_{21}\widehat{Q}-\lambda_{4}\widehat{\wp}_{21}+\left(\frac{\lambda_{5}}{2}+\frac{2\lambda_{0}\lambda_{6}}{\lambda_{1}}\right)\widehat{\wp}_{11}+\frac{4\lambda_{0}\lambda_{3}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{2}$
$-\frac{8\lambda_{0}\lambda_{4}}{\lambda_{1}^{2}}\widehat{\wp}_{21}\widehat{\wp}_{11}+\frac{4\lambda_{0}\lambda_{5}}{\lambda_{1}^{2}}\widehat{\wp}_{11}^{2}-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}\widehat{\wp}_{11}\widehat{Q}+\frac{\lambda_{1}\lambda_{6}}{4}=0,$ (3.40)
$7)\ \widehat{Q}_{22}-6\widehat{Q}^{2}-\lambda_{4}\widehat{Q}+3\lambda_{6}\widehat{\wp}_{11}-\left(\lambda_{5}-\frac{2\lambda_{0}\lambda_{6}}{\lambda_{1}}\right)\widehat{\wp}_{21}-\frac{8\lambda_{0}\lambda_{5}}{\lambda_{1}^{2}}\widehat{\wp}_{21}\widehat{\wp}_{11}$
$+\frac{16\lambda_{0}\lambda_{6}}{\lambda_{1}^{2}}\widehat{\wp}_{11}^{2}-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{2}\widehat{Q}+\left(\frac{\lambda_{2}\lambda_{6}}{2}-\frac{\lambda_{3}\lambda_{5}}{8}\right)=0.$ (3.41)
From Eq.(3.37), we naturally obtain $\widehat{Q}$ in the combination of Eq.(3.33). The characteristic feature of the $\lambda_{0}\neq 0$ hyperelliptic differential equations is the existence of the cubic terms ${-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{3},-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{2}\widehat{\wp}_{11},\cdots,-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{2}\widehat{Q}}$, in Eq.(3.30) $\sim$ Eq.(3.35). .
If we take the special case of $\lambda_{6}=0$, differential equations for the genus two Weierstrass type hyperelliptic functions does not contain cubic terms just as the differential equation of the genus one Weierstrass $\wp$ function. While the above differential equations for genus two Jacobi type hyperelliptic functions contain cubic terms just as the differential equation of the genus one Jacobi $\rm sn$} function. This is the reason why we call $\widehat{\wp}_{ij}$, $\widehat{Q}$ as the Jacobi type hyperelliptic functions.
In the more special case of $\lambda_{0}=0,\lambda_{6}=0$, we obtain ${\frac{\partial\widehat{Q}}{\partial u_{1}}=\frac{\partial\widehat{\wp}_{21}}{\partial u_{2}}}$, then we can put $\widehat{Q}=\widehat{\wp}_{22}$ in such special case. Then the correspondence of Eq.(3.33) becomes $\wp_{22}\rightarrow\widehat{\wp}_{11},\wp_{21}\rightarrow\widehat{\wp}_{21},\wp_{11}\rightarrow\widehat{\wp}_{22}$. In this special case, differential equations for the genus two Jacobi type hyperelliptic equations become
$1)\ \widehat{\wp}_{2222}-6\widehat{\wp}_{22}^{2}-\lambda_{4}\widehat{\wp}_{22}-\lambda_{5}\widehat{\wp}_{21}-\frac{\lambda_{3}\lambda_{5}}{8}=0,$ (3.42)
$2)\ \wp_{2221}-6\widehat{\wp}_{22}\wp_{21}-\lambda_{4}\widehat{\wp}_{21}+\frac{\lambda_{5}}{2}\widehat{\wp}_{11}=0,$ (3.43)
$3)\ \widehat{\wp}_{2211}-2\widehat{\wp}_{22}\widehat{\wp}_{11}-4\widehat{\wp}_{21}^{2}-\frac{\lambda_{3}}{2}\widehat{\wp}_{21}=0,$ (3.44)
$4)\ \widehat{\wp}_{2111}-6\wp_{21}\wp_{11}+\frac{\lambda_{1}}{2}\widehat{\wp}_{22}-\lambda_{2}\widehat{\wp}_{21}=0,$ (3.45)
$5)\ \widehat{\wp}_{1111}-6\widehat{\wp}_{11}^{2}-\lambda_{1}\widehat{\wp}_{21}-\lambda_{2}\widehat{\wp}_{11}-\frac{\lambda_{1}\lambda_{3}}{8}=0.$ (3.46)
Therefore, in the special case of $\lambda_{0}=0,\lambda_{6}=0$, the genus two Weierstrass type hyperelliptic functions $\{\wp_{22}.\wp_{21},\wp_{11}\}$ and the Jacobi type hyperelliptic functions $\{\widehat{\wp}_{22}.\widehat{\wp}_{21},\widehat{\wp}_{11}\}$ satisfy same set of differential equations. The relation between two solutions is given by
$\widehat{\wp}_{22}=-\frac{\lambda_{5}}{4}\frac{\wp_{11}}{\wp_{21}},\ \widehat{\wp}_{21}=\frac{\lambda_{1}\lambda_{5}}{16}\frac{1}{\wp_{21}},\ \widehat{\wp}_{11}=-\frac{\lambda_{1}}{4}\frac{\wp_{22}}{\wp_{21}},$ (3.47)
which corresponds the half-period relation
$(\wp(u+\omega_{3})-e_{3})=\frac{(e_{3}-e_{1})(e_{3}-e_{2})}{(\wp(u)-e_{3})},\quad{\rm sn}(u+3iK^{\prime})=\frac{1}{k\ {\rm sn}(u)},$
in genus one case. Eq.(3.47) is the special case of the half-period relation of the standard projective type II $\rm Sp(4,{\bf R})$} Lie group transformation
$\wp_{22}(u+\Omega)=\frac{a_{1}\wp_{22}(u)+a_{2}\wp_{21}(u)+a_{3}\wp_{11}(u)+a_{4}}{d_{1}\wp_{22}(u)+d_{2}\wp_{21}(u)+d_{3}\wp_{11}(u)+d_{4}},$ (3.48)
$\wp_{21}(u+\Omega)=\frac{b_{1}\wp_{22}(u)+b_{2}\wp_{21}(u)+b_{3}\wp_{11}(u)+b_{4}}{d_{1}\wp_{22}(u)+d_{2}\wp_{21}(u)+d_{3}\wp_{11}(u)+d_{4}},$ (3.49)
$\wp_{11}(u+\Omega)=\frac{c_{1}\wp_{22}(u)+c_{2}\wp_{21}(u)+c_{3}\wp_{11}(u)+c_{4}}{d_{1}\wp_{22}(u)+d_{2}\wp_{21}(u)+d_{3}\wp_{11}(u)+d_{4}},$ (3.50)
and
[…]
by putting $\lambda_{5}=4,\lambda_{1}=4$ [3].
4 Summary and Discussions
The static KdV equation has solutions of the Weierstrass $\wp$ function. The Weierstrass $\wp$ elliptic functions and the Jacobi $\rm sn$} elliptic functions form a family of elliptic functions. We have generalized the mKdV equation in such a way as the static mKdV differential equation and the static KdV differential equation have good correspondence. We have demonstrated that this generalized mKdV equation is an integrable equation. We have also demonstrated that there are four kinds of relations between solutions of the static KdV equation and the generalized static mKdV equation. Next, we generalized the relationship between $\wp$ and $\rm sn$} functions to the genus two case.
For the general genus two hyperelliptic curves, we obtained differential equations for the genus two Weierstrass type and Jacobi type hyperelliptic functions.
If we consider the special case of $\lambda_{6}=0,\lambda_{0}=0$, the differential equations of the genus two Weierstrass type and Jacobi type hyperelliptic functions become the same. Then the Weierstrass type and the Jacobi type hyperelliptic functions are linked as the half-period shift solutions of the same differential equations. In fact, we have confirmed that two hyperelliptic functions are linked by a special half-period $\rm Sp(4,{\bf R})$} Lie group transformation.

### v2

@@
 masahito.hayashi@oit.ac.jpshigemot@tezukayama-u.ac.jptsukioka@bukkyo-u.ac.jp
-We generalized the mKdV equation in order that the static equations include $\rm sn$} differential equation.
-As a result, a good correspondence was obtained between the KdV equation and the mKdV equation.
-For general genus two hyperelliptic curves, we obtained differential equations for Weierstrass type and Jacobi type hyperelliptic functions.
-Considering the special case of $\lambda_{6}=0,\lambda_{0}=0$, Weierstrass type and Jacobi type hyperelliptic functions are different solutions to the same hyperelliptic differential equations.
-Then these solutions are connected by the special $\rm Sp(4,{\bf R})$} Lie group transformation.
+We generalized the mKdV equation such that the static equations included the $\rm sn$} differential equation.
+As a result, a good correspondence was obtained between the KdV and mKdV equations.
+For the general genus two hyperelliptic curves, we obtained differential equations for the Weierstrass-type and Jacobi-type hyperelliptic functions.
+Considering the special case of $\lambda_{6}=0,\lambda_{0}=0$, the Weierstrass-type and Jacobi-type hyperelliptic functions are different solutions to the same hyperelliptic differential equations.
+These solutions are then connected via a special $\rm Sp(4,{\bf R})$} Lie group transformation.
 Keywords: generalized mKdV equation, genus two, Jacobi type, hyperelliptic differential equations
 1 Introduction
-The reason why soliton system can be solved exactly despite being nonlinear is that there is the Lie group structure.
-It is well known that the KdV equation has the ${\rm Sp(2,{\bf R})/Z_{2}}\cong{\rm SO(2,1)}$ Lie group structure [1, 2].
+The soliton system can be solved exactly despite being nonlinear because of the Lie group structure.
+It is well known that the KdV equation has an ${\rm Sp(2,{\bf R})/Z_{2}}\cong{\rm SO(2,1)}$ Lie group structure [1, 2].
 We have demonstrated that $\wp$ function, which is the special static solution of the KdV equation, has the ${\rm Sp(2,{\bf R})/Z_{2}}\cong{\rm SO(2,1)}$ Lie group structure [3], and the genus two hyperelliptic $\wp$ functions have the ${\rm Sp(4,{\bf R})/Z_{2}}\cong{\rm SO(3,2)}$ Lie group structure [3, 4, 5].
-Not in the general Riemann surfaces but in the hyperelliptic curves, we expect that multi-periodic functions, i.e., hyperelliptic functions have the Lie group structure.
-In this paper, we first generalized mKdV equation in order that the static differential equation includes $\rm sn$} differential equation.
-This gives a good correspondence between the KdV equation and the mKdV equation.
-For general genus two hyperelliptic curves, we obtain the conventional differential equations for Weierstrass type and Jacobi type hyperelliptic functions.
-If we consider the special case of $\lambda_{6}=0,\lambda_{0}=0$, differential equations for Weierstrass type and Jacobi type hyperelliptic functions are different solutions of the same hyperelliptic differential equations.
-We confirmed that two hyperelliptic functions are linked by a special half-period $\rm Sp(4,{\bf R})$} Lie group transformation.
+In the hyperelliptic curves, not in general Riemann surfaces, we expect that multi-periodic functions, that is, hyperelliptic functions, have a Lie group structure.
+In this study, we first generalized the mKdV equation such that the static differential equation included the $\rm sn$} differential equation.
+This gives a good correspondence between the KdV and mKdV equations.
+For the general genus two hyperelliptic curves, we obtain the conventional differential equations for the Weierstrass-type and Jacobi-type hyperelliptic functions.
+If we consider the special case of $\lambda_{6}=0,\lambda_{0}=0$, differential equations for the Weierstrass-type and Jacobi-type hyperelliptic functions are different solutions of the same hyperelliptic differential equations.
+We confirmed that the two hyperelliptic functions were linked by a special half-period $\rm Sp(4,{\bf R})$} Lie group transformation.
 2 Generalized mKdV equation
@@
 where we denote $\dfrac{\partial u(x,t)}{\partial t}=u_{t}(x,t)$, $\dfrac{\partial u(x,t)}{\partial x}=u_{x}(x,t)$, $\dfrac{\partial^{3}u(x,t)}{\partial x^{3}}=u_{xxx}(x,t)\rule$[-9.95845pt]{0.0pt}{25.60747pt} etc.
-Eq.(2.2) is generalized in order that the static equation includes differential equation for the $\rm sn$} function as follows
+Eq.(2.2) is generalized such that the static equation includes a differential equation for the $\rm sn$} function as follows
 $v_{t}(x,t)+v_{xxx}(x,t)-6v^{2}(x,t)v_{x}(x,t)+av_{x}(x,t)=0.$ (2.3)
@@
 The matrix $M$ in Eq.(2.21) has $\rm GL(2,{\bf R})$} structure.
-The integrability condition of Eqs.(2.1) and (2.2) is given by
+The integrability condition for Eq.(2.12) and (2.21) are given by
 $\left[\frac{\partial}{\partial x}-L,\frac{\partial}{\partial t}-M\right]=\frac{\partial L}{\partial t}-\frac{\partial M}{\partial x}+\left[L,M\right]=\left(\begin{array}[]{cc}0&D\\ -D&0\end{array}\right)=0,$ (2.27)
 $D=v_{t}(x,t)+v_{xxx}(x,t)-6v^{2}(x,t)v_{x}(x,t)+4\eta^{2}(b-1)v_{x}(x,t),$ (2.28)
-and $b$ is set to be $a=4\eta^{2}(b-1)$. Then the above integrability condition $D=0$ gives the generalized mKdV equation.
+and $b$ is set to be $a=4\eta^{2}(b-1)$. Then the above integrability condition $D=0$ yields the generalized mKdV equation.
 2.2 Various transformations between the static KdV and the static generalized mKdV equations
-For simplicity, we will denote $v(x,0)$ as $v(x)$. This type of abbreviation is used in other cases as well.
+For simplicity, we will denote $v(x,0)$ as $v(x)$. This type of abbreviation is also used in other cases.
 2.2.1 Miura transformation
@@
 $u_{xxx}(x)-6u(x)u_{x}(x)=0.$ (2.31)
-Of course, the time dependent KdV and generalized mKdV equations are linked by the time dependent Miura transformation $u(x,t)=v^{2}(x,t)+v_{x}(x,t)-a/6$ [8].
+The time-dependent KdV and generalized mKdV equations are linked by the time-dependent Miura transformation $u(x,t)=v^{2}(x,t)+v_{x}(x,t)-a/6$ [8].
 2.2.2 Square transformation
@@
 $\frac{d}{dx}\left(v_{x}^{2}(x)-v^{4}(x)+av^{2}(x)-\frac{1}{2}\right)=2v_{x}(x)\left(v_{xx}(x)-2v^{3}(x)+av(x)\right).$ (2.38)
-In this case, $v(x)$ not only satisfies the static generalized mKdV equation but also becomes the Jacobi $\rm sn$} function.
-This transformation is well-known that links the Weierstrass $\wp$ function and the Jacobi $\rm sn$} function in the form
+In this case, $v(x)$ not only satisfies the static generalized mKdV equation, but also becomes the Jacobi $\rm sn$} function.
+This transformation is well-known to link the Weierstrass $\wp$ function and Jacobi $\rm sn$} function in the form
 $\wp\left(\frac{u}{\sqrt{e_{1}-e_{3}}}\right)=e_{3}+\frac{(e_{1}-e_{3})}{{\rm sn}^{2}(u)}.$
@@
 $\frac{d}{dx}\left(v_{xx}(x)-2v^{3}(x)+av(x)-1\right)=v_{xxx}(x)-6v^{2}(x)v_{x}(x)+av_{x}(x).$
-In this case, $v(x)$ is a solution of the static generalized mKdV equation, but it is not that of the Jacobi $\rm sn$} function.
+In this case, $v(x)$ is a solution of the static generalized mKdV equation, but not that of the Jacobi $\rm sn$} function.
 2.2.5 A relationship between transformations
@@
 $u(x)=2{v_{1}^{2}(x)}-\frac{2a}{3}=\frac{1}{{v_{2}^{2}(x)}}-\frac{2a}{3},$ (2.41)
-where $v_{1}(x)$ and $v_{2}(x)$ satisfies the same $\rm sn$} type differential equation.
-In order that Eq.(2.40) becomes the standard $\rm sn$} differential equation, we put $k=\sqrt{2}$, $a=(1+k^{2})/2=3/2$, and we obtain [3]
+where $v_{1}(x)$ and $v_{2}(x)$ satisfy the same $\rm sn$}-type differential equation.
+In order that Eq.(2.40) becomes the standard $\rm sn$} differential equation, we set $k=\sqrt{2}$, $a=(1+k^{2})/2=3/2$, and obtain
 $v_{1}(x)={\rm sn}(x/\sqrt{2})~~\mathrm{and}~~v_{2}(x)=\widehat{{\rm sn}}(x/\sqrt{2})=\frac{1}{\sqrt{2}\ {\rm sn}(x/\sqrt{2})}=\frac{1}{\sqrt{2}\ v_{1}(x)}.$ (2.42)
@@
 ${\rm sn}_{zz}(z)=-\left(1+k^{2}\right){\rm sn}(z)+2k^{2}{\rm sn}^{3}(z).$ (2.43)
-On the other hand, ${\widehat{{\rm sn}}(z)=\frac{1}{k\,{\rm sn}(z)}}$ satisfies the same differential equation
+On the other hand, ${\widehat{{\rm sn}}(z)=\frac{1}{k\,{\rm sn}(z)}}$ satisfies the same differential equation [3]
 $\widehat{{\rm sn}}_{zz}(z)=-\left(1+k^{2}\right)\widehat{{\rm sn}}(z)+2k^{2}\widehat{{\rm sn}}^{3}(z).$ (2.44)
@@
 Though $\Re_{22}$ and $\Re_{21}$ are symmetric functions, but they are not the fundamental symmetric functions.
-Then any symmetric functions cannot be expressed in the closed form by $\Re_{22}$ and $\Re_{21}$.
-So far, closed differential equations for the genus two hyperelliptic functions are known only for the $\lambda_{6}=0$ case [10].
+Therefore, it is not always possible to express any symmetric function in a closed form using $\Re_{22}$ and $\Re_{21}$.
+Until now, closed differential equations for the genus two hyperelliptic functions are known only for the $\lambda_{6}=0$ case [10].
 3.1 Genus two Weierstrass type hyperelliptic differential equations
@@
 $\wp_{22}(u_{1},u_{2})=\frac{\lambda_{5}}{4}(x_{1}+x_{2}),\ \wp_{21}(u_{1},u_{2})=-\frac{\lambda_{5}}{4}x_{1}x_{2},\ Q(u_{1},u_{2})=\frac{F(x_{1},x_{2};\left\{\lambda_{j}\right\})-2y_{1}y_{2}}{4(x_{1}-x_{2})^{2}}.$ (3.9)
-We call the above as the Weierstrass type hyperelliptic functions.
+We refer to this as Weierstrass-type hyperelliptic functions.
 Here $\wp_{22}$ and $\wp_{21}$ are fundamental symmetric function.
-Then any symmetric quantities can be expressed by $\wp_{22}$ and $\wp_{21}$.
-Thus we can obtain the hyperelliptic differential equations in the closed forms.
+Then any symmetric quantity can be expressed by $\wp_{22}$ and $\wp_{21}$.
+Thus, we can obtain hyperelliptic differential equations in closed forms.
 Though we obtain the relation
 $\frac{\partial\wp_{22}}{\partial u_{1}}=\frac{\partial\wp_{21}}{\partial u_{2}},\quad\frac{\partial\wp_{21}}{\partial u_{1}}\neq\frac{\partial Q}{\partial u_{2}},$ (3.10)
-the condition ${\frac{\partial\wp_{22}}{\partial u_{1}}=\frac{\partial\wp_{21}}{\partial u_{2}}}$ alone is sufficient to guarantee the integrability condition, i.e.
-the existence of $\sigma(u_{1},u_{2})$ function as a potential.
+the condition ${\frac{\partial\wp_{22}}{\partial u_{1}}=\frac{\partial\wp_{21}}{\partial u_{2}}}$ alone is sufficient to guarantee the integrability condition, that is, the existence of $\sigma(u_{1},u_{2})$ function as a potential.
 If $\wp_{22}$ and $\wp_{21}$ are given as functions of $\sigma$ function, $Q$ can in principle be given as a function of $\sigma$ function via the generalized Kummer surface relation.
-We will give the generalized Kummer surface relation later.
+We will provide a generalized Kummer surface relation later.
 In order to obtain the differential equations for genus two Weierstrass type hyperelliptic functions, we first calculate $\wp_{2222}$, $\wp_{2221}$, $\wp_{2211}$, $\wp_{2111}$, $Q_{11}$, $Q_{21}$ and $Q_{22}$.
 Because these quantities are symmetric functions of $x_{1}$ and $x_{2}$, these are expressed by $\wp_{22}$ and $\wp_{21}$.
-If these quantities contain symmetric function of the type $y_{1}y_{2}$, the quantity $y_{1}y_{2}$ is expressed by $Q$ and/or $Q^{2}$ by using ${y_{1}y_{2}=-2(x_{1}-x_{2})^{2}Q+\frac{1}{2}F(x_{1},x_{2})}$,
+If these quantities contain symmetric function of the type $y_{1}y_{2}$, the quantity $y_{1}y_{2}$ is expressed by $Q$ and/or $Q^{2}$ by using ${y_{1}y_{2}=-2(x_{1}-x_{2})^{2}Q+\frac{1}{2}F(x_{1},x_{2})}$.
 Then we obtain the differential equations of the genus two Weierstrass type hyperelliptic functions of the form
@@
 $-\frac{20\lambda_{6}}{\lambda_{5}}\wp_{21}Q+\frac{\lambda_{1}\lambda_{6}}{\lambda_{5}}\wp_{22}+\left(-\frac{2\lambda_{2}\lambda_{6}}{\lambda_{5}}-\frac{\lambda_{3}}{2}\right)\wp_{21}+\frac{\lambda_{0}\lambda_{6}}{2}=0.$ (3.17)
-From Eq.(3.12), we naturally obtain $Q$ in the combination of Eq.(3.9).
+From Eq.(3.12), $Q$ naturally appears as a combination of the terms in Eq.(3.9).
 The characteristic feature of $\lambda_{6}\neq 0$ hyperelliptic differential equations is the existence of the cubic term, that is, terms ${-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{3},-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{2}\wp_{21},\cdots,-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{2}Q}$, in Eq.(3.11) $\sim$ Eq.(3.17).
-Here we give the generalized Kummer surface relation.
+Here we present a generalized Kummer surface relation.
 For $\wp_{22}$, $\wp_{21}$ and $Q$, only two are independent.
-Then there exist one relation, which is called the generalized Kummer surface relation.
-By using the relation ${y_{1}y_{2}=-2(x_{1}-x_{2})^{2}Q+\frac{1}{2}F(x_{1},x_{2})}$, we obtain
+Then, there exists one relation, called the generalized Kummer surface relation.
+Using the relation ${y_{1}y_{2}=-2(x_{1}-x_{2})^{2}Q+\frac{1}{2}F(x_{1},x_{2})}$, we obtain
 ${y_{1}^{2}y_{2}^{2}=4(x_{1}-x_{2})^{4}Q^{2}-2(x_{1}-x_{2})^{2}QF(x_{1},x_{2})+\frac{1}{4}F(x_{1},x_{2})^{2}}.$
 Using Eq.(3.1), this constraint is expressed by $\wp_{22}$, $\wp_{21}$ and $Q$.
-The explicit expression is given as follows.
-We denote $X=\wp_{22}$, $Y=\wp_{21}$, $Z=Q$, and we obtain the following generalized Kummer surface relation of the form
+The explicit expression is as follows.
+We denote $X=\wp_{22}$, $Y=\wp_{21}$, $Z=Q$, and obtain the following generalized Kummer surface relation of the form
 $K_{2}=K_{1}+\frac{4\lambda_{6}}{\lambda_{5}^{2}}\Big(-\lambda_{5}^{2}Y^{2}-8\lambda_{0}\lambda_{5}X^{2}Y+4\lambda_{1}\lambda_{5}XY^{2}$
@@
 If we consider the special case of $\lambda_{6}=0$, we obtain ${\frac{\partial\wp_{21}}{\partial u_{1}}=\frac{\partial Q}{\partial u_{2}}}$, then we can put $Q=\wp_{11}$.
-Thus we obtain the conventional differential equations for Weierstrass type hyperelliptic functions [11].
-If we consider the more special case of $\lambda_{6}=0,\lambda_{0}=0$, differential equations for Weierstrass type hyperelliptic differential functions become
+Thus, we obtain conventional differential equations for Weierstrass-type hyperelliptic functions [11].
+If we consider the more special case of $\lambda_{6}=0,\lambda_{0}=0$, the differential equations for the Weierstrass-type hyperelliptic differential functions become
 $1)\ \wp_{2222}-6\wp_{22}^{2}-\lambda_{4}\wp_{22}-\lambda_{5}\wp_{21}-\frac{\lambda_{3}\lambda_{5}}{8}=0,$ (3.24)
@@
 $5)\ \wp_{1111}-6\wp_{11}^{2}-\lambda_{1}\wp_{21}-\lambda_{2}\wp_{11}-\frac{\lambda_{1}\lambda_{3}}{8}=0.$ (3.28)
-The condition $\lambda_{0}=0$ is not the special restriction, because we can always put $\lambda_{0}=0$ by the constant shift of $x_{i}$ in the hyperelliptic curve.
+The condition $\lambda_{0}=0$ is not a special restriction, because we can always put $\lambda_{0}=0$ by the constant shift of $x_{i}$ in the hyperelliptic curve.
 3.2 Genus two Jacobi type hyperelliptic differential equations
-Next, we try to obtain differential equations for the genus two Jacobi type hyperelliptic functions by the following dual transformation. Under the dual transformation
+Next, we try to obtain differential equations for the genus two Jacobi-type hyperelliptic functions using the following dual transformation. Under a dual transformation
 $x_{i}\rightarrow\frac{1}{x_{i}},\quad y_{i}\rightarrow\frac{y_{i}}{x_{i}^{3}},\quad\lambda_{j}\leftrightarrow\lambda_{6-j},\ (j=0,1,\cdots,6),$
@@
 $+\frac{16\lambda_{0}\lambda_{6}}{\lambda_{1}^{2}}\widehat{\wp}_{11}^{2}-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{2}\widehat{Q}+\left(\frac{\lambda_{2}\lambda_{6}}{2}-\frac{\lambda_{3}\lambda_{5}}{8}\right)=0.$ (3.41)
-From Eq.(3.37), we naturally obtain $\widehat{Q}$ in the combination of Eq.(3.33).
+From Eq.(3.37), $\widehat{Q}$ naturally appears as a combination of the terms in Eq.(3.33).
 The characteristic feature of the $\lambda_{0}\neq 0$ hyperelliptic differential equations is the existence of the cubic terms ${-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{3},-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{2}\widehat{\wp}_{11},\cdots,-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{2}\widehat{Q}}$, in Eq.(3.35) $\sim$ Eq.(3.41).
-.
 If we take the special case of $\lambda_{6}=0$, differential equations for the genus two Weierstrass type hyperelliptic functions does not contain cubic terms just as the differential equation of the genus one Weierstrass $\wp$ function.
-While the above differential equations for genus two Jacobi type hyperelliptic functions contain cubic terms just as the differential equation of the genus one Jacobi $\rm sn$} function.
-This is the reason why we call $\widehat{\wp}_{ij}$, $\widehat{Q}$ as the Jacobi type hyperelliptic functions.
-In the more special case of $\lambda_{0}=0,\lambda_{6}=0$, we obtain ${\frac{\partial\widehat{Q}}{\partial u_{1}}=\frac{\partial\widehat{\wp}_{21}}{\partial u_{2}}}$, then we can put $\widehat{Q}=\widehat{\wp}_{22}$ in such special case.
-Then the correspondence of Eq.(3.33) becomes $\wp_{22}\rightarrow\widehat{\wp}_{11},\wp_{21}\rightarrow\widehat{\wp}_{21},\wp_{11}\rightarrow\widehat{\wp}_{22}$.
+The above differential equations for genus two Jacobi-type hyperelliptic functions contain cubic terms just as the differential equation of the genus one Jacobi $\rm sn$} function.
+This is why we call $\widehat{\wp}_{ij}$, $\widehat{Q}$ Jacobi-type hyperelliptic functions.
+In the special case of $\lambda_{0}=0,\lambda_{6}=0$, we obtain ${\frac{\partial\widehat{Q}}{\partial u_{1}}=\frac{\partial\widehat{\wp}_{21}}{\partial u_{2}}}$, then we can put $\widehat{Q}=\widehat{\wp}_{22}$ in such a special case.
+Therefore, the correspondence in Eq.(3.33) is as follows: $\wp_{22}\rightarrow\widehat{\wp}_{11},\wp_{21}\rightarrow\widehat{\wp}_{21},\wp_{11}\rightarrow\widehat{\wp}_{22}$.
 In this special case, differential equations for the genus two Jacobi type hyperelliptic equations become
@@
 $(\wp(u+\omega_{3})-e_{3})=\frac{(e_{3}-e_{1})(e_{3}-e_{2})}{(\wp(u)-e_{3})},\quad{\rm sn}(u+3iK^{\prime})=\frac{1}{k\ {\rm sn}(u)},$
-in genus one case. Eq.(3.47) is the special case of the half-period relation of the standard projective type II $\rm Sp(4,{\bf R})$} Lie group transformation
+in genus one case. Eq.(3.47) is a special case of the half-period relation of the standard projective type II $\rm Sp(4,{\bf R})$} Lie group transformation
 $\wp_{22}(u+\Omega)=\frac{a_{1}\wp_{22}(u)+a_{2}\wp_{21}(u)+a_{3}\wp_{11}(u)+a_{4}}{d_{1}\wp_{22}(u)+d_{2}\wp_{21}(u)+d_{3}\wp_{11}(u)+d_{4}},$ (3.48)
@@
 The static KdV equation has solutions of the Weierstrass $\wp$ function.
-The Weierstrass $\wp$ elliptic functions and the Jacobi $\rm sn$} elliptic functions form a family of elliptic functions.
-We have generalized the mKdV equation in such a way as the static mKdV differential equation and the static KdV differential equation have good correspondence.
-We have demonstrated that this generalized mKdV equation is an integrable equation.
-We have also demonstrated that there are four kinds of relations between solutions of the static KdV equation and the generalized static mKdV equation.
+The Weierstrass $\wp$ elliptic function and Jacobi $\rm sn$} elliptic functions form a family of elliptic functions.
+We have generalized the mKdV equation in such a way that the static mKdV differential equation and static KdV differential equation have good correspondence.
+We demonstrated that this generalized mKdV equation is integrable.
+We have also demonstrated that there are four kinds of relations between the solutions of the static KdV equation and the generalized static mKdV equation.
 Next, we generalized the relationship between $\wp$ and $\rm sn$} functions to the genus two case.
-For the general genus two hyperelliptic curves, we obtained differential equations for the genus two Weierstrass type and Jacobi type hyperelliptic functions.
-If we consider the special case of $\lambda_{6}=0,\lambda_{0}=0$, the differential equations of the genus two Weierstrass type and Jacobi type hyperelliptic functions become the same.
-Then the Weierstrass type and the Jacobi type hyperelliptic functions are linked as the half-period shift solutions of the same differential equations.
-In fact, we have confirmed that two hyperelliptic functions are linked by a special half-period $\rm Sp(4,{\bf R})$} Lie group transformation.
+For the general genus two hyperelliptic curves, we obtained differential equations for the genus two Weierstrass-type and Jacobi-type hyperelliptic functions.
+If we consider the special case of $\lambda_{6}=0,\lambda_{0}=0$, the differential equations of the genus two Weierstrass-type and Jacobi-type hyperelliptic functions become the same.
+Then the Weierstrass-type and the Jacobi-type hyperelliptic functions are linked as half-period shift solutions of the same differential equations.
+In fact, we confirmed that two hyperelliptic functions are linked by a special half-period $\rm Sp(4,{\bf R})$} Lie group transformation.